# Weak-Coupling Limit of the Lattice Nonlinear Schrödinger Integral Equation

### 10 Mar 2026 v1

@@ A.2 Resolvent amplification and the origin of logarithmic growth
The tail sum is controlled by the gap between $\lambda_{0}$ and $\lambda_{1}$. Since the spectral gaps $2\pi-\lambda_{n}$ for $n\geq 1$ remain bounded below by $\Delta_{1}(Q)=2\pi-\lambda_{1}(Q)>\Delta_{0}(Q)$ (and in fact $\Delta_{1}/\Delta_{0}\to\infty$ numerically, cf. Table 3 in the main text), the tail satisfies
@@
using Cauchy–Schwarz. By the Mercer expansion (119) at $\xi=\eta=0$, $\sum_{n}\lambda_{n}\,\phi_{n}(0)^{2}=K(0)=2$, so $\sum_{n}\phi_{n}(0)^{2}\leq 2/\lambda_{\min}$, which is finite. Since $\|g\|_{L^{2}}^{2}=\int_{-Q}^{Q}4/(1+\xi^{2})^{2}\,d\xi\leq 2\pi$ and $\Delta_{1}(Q)$ remains bounded away from zero (or grows slower than $\Delta_{0}^{-1}$), the tail is $\mathcal{O}(1)$. ∎
To extract the precise logarithmic behaviour from (133), we need the projections of the driving term and the eigenfunction onto the leading mode. As $Q\to\infty$, the leading eigenfunction approaches a normalised constant: $\phi_{0}(\xi)\to(2Q)^{-1/2}$, giving
$\langle\phi_{0},g\rangle\;\approx\;(2Q)^{-1/2}\int_{-Q}^{Q}\frac{2}{1+\xi^{2}}\,d\xi\;=\;(2Q)^{-1/2}\,(2\pi-4/Q+\mathcal{O}(Q^{-3}))\,$, (136)
@@ A.3 Fredholm determinant and eigenvalue product formula
The Szegő–Widom theory [13] provides the asymptotics of $\mathcal{F}(Q)$, taking into account the zero of the symbol $\sigma(p)=1-e^{-|p|}$ at $p=0$:
@@
where $\pi/3$ is the Szegő coefficient (Proposition 7.1 in the main text) and $\alpha_{\mathrm{FH}}=1/4$ is the Fisher–Hartwig exponent arising from the simple zero of $\sigma$ at $p=0$.
@@
$\frac{1}{2Q}\,\log\mathcal{F}(Q)\;\to\;\frac{1}{2\pi}\int_{-\infty}^{\infty}\log\sigma(p)\,dp\;=\;\frac{1}{\pi}\int_{0}^{\infty}\log(1-e^{-p})\,dp\;=\;-\frac{\pi}{6}\,$, (141)
where we used $\int_{0}^{\infty}\log(1-e^{-p})\,dp=-\operatorname{Li}_{2}(1)=-\pi^{2}/6$ (Proposition 7.1). Multiplying by $2Q$ gives the leading term $\pi Q/3$.
@@
are controlled by the complex zeros of $\sigma(p)=1-e^{-p}$ at $p_{n}=2\pi in$ ($n\in\mathbb{Z}\setminus\{0\}$), located at distance $2\pi$ from the real axis. This identifies the instanton action $A=2\pi$, as discussed in Section 7 of the main text.
@@ A.4 Numerical computation of eigenvalues
Table 4: Eigenvalues $\lambda_{n}(Q)$ of $\mathcal{K}_{Q}$ for various $Q$, computed by NN-point Gauss–Legendre discretisation with $N=10Q+400$. All eigenvalues lie below $2\pi=6.2832\ldots$ and converge to this value from below as $Q\to\infty$.
@@ Appendix B The total density expansion
where $b\approx-0.2173$ numerically.
We now derive $a=1/(2\pi)$ in the expansion (55). From the duality (53), decomposing $f(\xi)=f(0)+[f(\xi)-f(0)]$:
@@
and using $1-\frac{2\arctan Q}{\pi}=\frac{2}{\pi Q}+\mathcal{O}(Q^{-3})$:
@@
This key formula converts the problem to computing $R(Q)$.
@@
$R(Q)=(2-\pi)+\frac{2a\log Q+(2b+1)}{Q}+\cdots\,.$ (150)
In the inner region ($|\xi|\lesssim\mathcal{O}(1)$), both $f$ and $\tilde{\rho}$ are amplified by the same near-critical resolvent. The spectral projection gives $f(\xi)=2Q\,\tilde{\rho}(\xi)+f_{\perp}(\xi)$, where $f_{\perp}$ is bounded. Using the established inner solution $\tilde{\rho}(\xi)=\frac{1}{\pi}[\log(2Q)+\gamma_{E}-\operatorname{Re}\psi(1+i\xi)-\gamma_{E}]+\cdots\,$:
$f(\xi)=\frac{2Q}{\pi}\left[\log(2Q)-\operatorname{Re}\psi(1+i\xi)\right]+f_{\perp}(\xi)+\cdots\,.$ (151)
@@
$f(0)=\frac{2Q}{\pi}\left[\log(2Q)+\gamma_{E}\right]+f_{\perp}(0)+\cdots\,.$ (152)
The difference $f(\xi)-f(0)$ in the inner region is:
$f(\xi)-f(0)=\frac{2Q}{\pi}\left[-\operatorname{Re}\psi(1+i\xi)-\gamma_{E}\right]+[f_{\perp}(\xi)-f_{\perp}(0)]+\cdots\,.$ (153)
@@
where we applied the digamma integral identity  (56); the factor of 2 accounts for both halves of the symmetric integral), and $R_{\perp}(Q):=\int[f_{\perp}(\xi)-f_{\perp}(0)]/(1+\xi^{2})\,d\xi$.
@@
$D=\frac{Q[\pi-2Q+R_{\perp}(Q)]}{2}-\frac{1}{2}+\mathcal{O}(Q^{-1})=\frac{Q\pi}{2}-Q^{2}+\frac{Q\,R_{\perp}(Q)}{2}-\frac{1}{2}\,.$ (155)
@@
The expansion (156) is physically expected: $f_{\perp}$ is the non-critical component of the Love solution, which equals $f-2Q\tilde{\rho}$ and grows as $Q$ in the outer region (the Fermi sea contribution). In the outer region, $f_{\perp}(\xi)\approx f_{\mathrm{bulk}}-2Q\cdot\tfrac{1}{2}=f_{\mathrm{bulk}}-Q$. Since $f_{\mathrm{bulk}}\approx Q$, we get $f_{\perp}\sim 0$ in the deep outer region, but $f_{\perp}(0)\approx f(0)-2Q\tilde{\rho}(0)=(1+Q+a\log Q+b)-\frac{2Q}{\pi}(\log(2Q)+\gamma_{E})$, which is $\mathcal{O}(\log Q)$ after the $O(Q)$ terms cancel. The Lorentzian-weighted integral of $f_{\perp}-f_{\perp}(0)$ thus picks up a contribution $\sim\log Q/Q$ from the slowly varying mismatch, producing the logarithmic coefficient.
@@
$D=Q+\frac{(2-\pi)Q}{2}+\frac{Q\pi}{2}-Q+a\log Q+\ldots=Q+a\log Q+\ldots\,$, (157)
confirming the structure. The coefficient $a$ is determined by the $O(1/Q)$ correction to the Lorentzian integral in (61) on the finite domain:
$\int_{0}^{Q}\frac{1-\cos(\xi t)}{1+\xi^{2}}\,d\xi=\frac{\pi}{2}(1-e^{-t})-\frac{1}{Q}h(t)+\mathcal{O}(Q^{-2})\,$, (158)
@@
The key mechanism is: the digamma integral identity (56) controls the leading $O(Q)$ part of $R(Q)$ (producing the cancellation that keeps $D\sim Q$ rather than $Q^{2}$), while the subleading correction, involving the interplay between the finite-domain boundary at $\xi=\pm Q$ and the Bose–Einstein mode structure, produces the coefficient $a=1/(2\pi)$.
@@ Appendix C Wiener–Hopf factorisation
where $K_{+}$ ($K_{-}$) is analytic and nonzero in the upper (lower) half of the complex $p$-plane. The derivation proceeds in five steps: we first express $\Sigma$ as a product of three elementary factors, then split each factor between the two half-planes, and finally verify the result.
Step 1. The gamma-function identity. The Euler reflection formula states
@@
The linear zero of $\Sigma$ is thus distributed as a square-root singularity between the two factors.
@@
$G_{+}(z)=\frac{K_{+}(z)}{(-iz)^{1/2}}=\frac{1}{\Gamma\!\left(1-\frac{iz}{2\pi}\right)}\,\exp\!\left[-\frac{iz}{2\pi}\log(-iz)\right],\qquad G_{-}(z)=\frac{K_{-}(z)}{(iz)^{1/2}}\,.$ (181)
Since $\Gamma(1)=1$ and the exponential tends to unity at $z=0$:
@@ Appendix D Preliminary numerical extraction of perturbative coefficients
We solve the integral equation at $M=60$ values $Q_{j}\in[20,\,500]$ using $N(Q)=10Q+400$ quadrature points, subtract the known leading terms, and fit to the truncated ansatz $R_{0}(Q)=\sum_{n=1}^{N_{\max}}(a_{n0}+a_{n1}\log Q)/Q^{n}$ by least squares, progressively increasing $N_{\max}$.
The design matrix has condition number growing explosively with $N_{\max}$: only the first 9–10 SVD modes carry signal, corresponding to roughly 5 reliable coefficient pairs $(a_{n0},a_{n1})$. The stable coefficients ($n=1,\ldots,5$) exhibit an alternating sign pattern (+,,+,,+)(+,-,+,-,+), consistent with a Borel singularity on the positive real axis.
@@
The ratio test $r_{n}/n=a_{(n+1)0}/(n\,a_{n0})$ should converge to $1/(2\pi)\approx 0.159$. The computed ratios (using $n\leq 5$) give $r_{n}/n\approx 0.6$$1.0$, still far from the predicted value—this slow convergence is expected since the asymptotic regime requires $n\gg 2\pi\approx 6.3$, and with only 5 reliable coefficients we are probing the pre-asymptotic regime.
A more direct test is to subtract the $N$-term perturbative approximation and look for the exponential tail. However, a quantitative analysis reveals that clean detection of the instanton is fundamentally impossible in double precision: at any $Q$, the perturbative remainder dominates $e^{-2\pi Q}$ by 9 or more orders of magnitude. The machine-epsilon threshold $e^{-2\pi Q}\approx 10^{-16}$ occurs at $Q\approx 5.9$, but at $Q=5.9$ the perturbative remainder after subtracting 4 terms is still $\sim\mathcal{O}(1)$. Extending to 15–20 reliable coefficients via extended-precision arithmetic would open a detection window.

### 22 Mar 2026 v2

@@ A.2 Resolvent amplification and the origin of logarithmic growth
The tail sum is controlled by the gap between $\lambda_{0}$ and $\lambda_{1}$. Since the spectral gaps $2\pi-\lambda_{n}$ for $n\geq 1$ remain bounded below by $\Delta_{1}(Q)=2\pi-\lambda_{1}(Q)>\Delta_{0}(Q)$ (and in fact $\Delta_{1}/\Delta_{0}\to\infty$ numerically), the tail satisfies
@@
using Cauchy–Schwarz. By the Mercer expansion (143) at $\xi=\eta=0$, $\sum_{n}\lambda_{n}\,\phi_{n}(0)^{2}=K(0)=2$, so $\sum_{n}\phi_{n}(0)^{2}\leq 2/\lambda_{\min}$, which is finite. Since
$\|g\|_{L^{2}}^{2}=\int_{-Q}^{Q}\frac{4}{(1+\xi^{2})^{2}}\,d\xi\leq 2\pi$ (161)
and $\Delta_{1}(Q)$ remains bounded away from zero (or grows slower than $\Delta_{0}^{-1}$), the tail is $\mathcal{O}(1)$. ∎
To extract the precise logarithmic behaviour from (158), we need the projections of the driving term and the eigenfunction onto the leading mode. As $Q\to\infty$, the leading eigenfunction approaches a normalised constant: $\phi_{0}(\xi)\to(2Q)^{-1/2}$, resulting in
$\langle\phi_{0},g\rangle\approx\sqrt{\frac{1}{2Q}}\int_{-Q}^{Q}\frac{2}{1+\xi^{2}}\,d\xi=\sqrt{\frac{1}{2Q}}\left(2\pi-\frac{4}{Q}\right)\,$, (162)
@@ A.3 Fredholm determinant and eigenvalue product formula
The Szegő–Widom theory [13] provides the asymptotics of $\mathcal{F}(Q)$, taking into account the zero of $\sigma(p)$ at $p=0$:
@@
$\log\mathcal{F}(Q)\;=\;-\frac{\pi Q}{3}\;+\;\alpha_{\mathrm{FH}}\,\log Q\;+\;\beta_{\mathrm{FH}}\;+\;\mathcal{O}\left(Q^{-1}\right)\,$, (166)
where $\pi/3$ is the Szegő coefficient and $\alpha_{\mathrm{FH}}=1/4$ is the Fisher–Hartwig exponent arising from the simple zero of $\sigma$ at $p=0$.
@@
$\frac{1}{2Q}\,\log\mathcal{F}(Q)\;\to\;\frac{1}{2\pi}\int_{-\infty}^{\infty}\log\sigma(p)\,dp\;=\;\frac{1}{\pi}\int_{0}^{\infty}\log(1-e^{-p})\,dp\;=\;-\frac{\pi}{6}\,.$ (167)
Multiplying by $2Q$ provides the leading term $\pi Q/3$.
@@
are controlled by the complex zeros of $\sigma(p)=1-e^{-p}$ at $p_{n}=2\pi in$ ($n\in\mathbb{Z}\setminus\{0\}$), located at distance $2\pi$ from the real axis. This identifies the instanton action $A=2\pi$, as discussed in Section 8.
@@ A.4 Numerical computation of eigenvalues
Table 4: Eigenvalues $\lambda_{n}(Q)$ of $\mathcal{K}_{Q}$ for various $Q$, computed by NN-point Gauss–Legendre discretisation with $N=10Q+400$. All eigenvalues lie below $2\pi$ and converge to this value from below as $Q\to\infty$.
@@ Appendix B The total density expansion
where $b\approx-0.2173$.
We start by deriving (74) in the expansion (57). From the duality (55), decomposing $f(\xi)=f(0)+[f(\xi)-f(0)]$:
@@
and using
$1-\frac{2\arctan Q}{\pi}=\frac{2}{\pi Q}+\mathcal{O}(Q^{-3})\,$, (175)
we have
@@
This is a key formula that converts the problem to computing $R(Q)$.
@@
$R(Q)=(2-\pi)+\frac{2a\log Q+(2b+1)}{Q}\,.$ (177)
In the inner region ($|\xi|\lesssim\mathcal{O}(1)$), both $f$ and $\tilde{\rho}$ are amplified by the same near-critical resolvent. The spectral projection gives $f(\xi)=2Q\,\tilde{\rho}(\xi)+f_{\perp}(\xi)$, where $f_{\perp}$ is bounded. Using the established inner solution (47), we have
$f(\xi)=\frac{2Q}{\pi}\left[\log(2Q)-\operatorname{Re}\psi(1+i\xi)\right]+f_{\perp}(\xi)\,.$ (178)
@@
$f(0)=\frac{2Q}{\pi}\left[\log(2Q)+\gamma_{E}\right]+f_{\perp}(0)\,.$ (179)
The difference $f(\xi)-f(0)$ in the inner region is then
$f(\xi)-f(0)=\frac{2Q}{\pi}\left[-\operatorname{Re}\psi(1+i\xi)-\gamma_{E}\right]+[f_{\perp}(\xi)-f_{\perp}(0)]\,.$ (180)
@@
where we applied the digamma integral identity  (58) (the factor of 2 accounts for both halves of the symmetric integral), and
$R_{\perp}(Q):=\int_{-\infty}^{\infty}\frac{f_{\perp}(\xi)-f_{\perp}(0)}{1+\xi^{2}}\,d\xi\,.$ (182)
@@
$D=\frac{1}{2}Q[\pi-2Q+R_{\perp}(Q)]-\frac{1}{2}+\mathcal{O}(Q^{-1})=\frac{Q\pi}{2}-Q^{2}+\frac{Q\,R_{\perp}(Q)}{2}-\frac{1}{2}\,.$ (183)
@@
The expansion (184) is physically expected: $f_{\perp}$ is the non-critical component of the Love solution, which equals $f-2Q\tilde{\rho}$ and grows as $Q$ in the outer region (the Fermi sea contribution). In the outer region, $f_{\perp}(\xi)\approx f_{\mathrm{bulk}}-2Q\cdot 1/2$. Since $f_{\mathrm{bulk}}\approx Q$, we get $f_{\perp}\sim 0$ in the deep outer region, but
$f_{\perp}(0)\approx f(0)-2Q\tilde{\rho}(0)=(1+Q+a\log Q+b)-\frac{2Q}{\pi}(\log(2Q)+\gamma_{E})\,$, (185)
which is $\mathcal{O}(\log Q)$ after the $\mathcal{O}(Q)$ terms cancel. The Lorentzian-weighted integral of $f_{\perp}-f_{\perp}(0)$ thus picks up a contribution $\sim\log Q/Q$ from the slowly varying mismatch, producing the logarithmic coefficient.
@@
$D=Q+\frac{(2-\pi)Q}{2}+\frac{Q\pi}{2}-Q+a\log Q=Q+a\log Q\,$, (186)
confirming the structure. The coefficient $a$ is determined by the $\mathcal{O}(1/Q)$ correction to the Lorentzian integral in (63) on the finite domain:
$\int_{0}^{Q}\frac{1-\cos(\xi t)}{1+\xi^{2}}\,d\xi=\frac{\pi}{2}(1-e^{-t})-\frac{1}{Q}h(t)+\mathcal{O}\left(Q^{-2}\right)\,$, (187)
@@
The key mechanism is: the digamma integral identity (58) controls the leading $\mathcal{O}(Q)$ part of $R(Q)$ (producing the cancellation that keeps $D\sim Q$ rather than $Q^{2}$), while the subleading correction, involving the interplay between the finite-domain boundary at $\xi=\pm Q$ and the Bose–Einstein mode structure, produces the coefficient $a=1/(2\pi)$ (74).
The profile integral (72) provides a complementary perspective. The inner–outer composite $\tilde{\rho}_{c}=1/2-\Phi(\xi)/\pi$ contributes
$2\int_{0}^{Q}\left[\tilde{\rho}_{c}-\tfrac{1}{2}\right]d\xi=-\frac{2}{\pi}\int_{0}^{Q}\Phi(\xi)\,d\xi=\frac{1}{2}+\mathcal{O}(1/Q)$ (189)
to D(Q)QD(Q)-Q. Therefore the inner and outer regions together contribute exactly $1/2$, and the entire non-trivial content of the expansion—the $1/(2\pi)\cdot\log Q$ term and the constant bb—is generated by the edge boundary layers.
@@ Appendix C Wiener–Hopf factorisation
where $K_{+}$ ($K_{-}$) is analytic and nonzero in the upper (lower) half of the complex $p$-plane. The derivation proceeds in five steps: we first express $\Sigma(p)$ as a product of three elementary factors, then split each factor between the two half-planes, and finally verify the result.
Step 1. Gamma-function identity. The Euler reflection formula states
@@
The linear zero of $\Sigma(p)$ is thus distributed as a square-root singularity between the two factors.
@@
$G_{+}(z)$ $=\frac{K_{+}(z)}{\sqrt{-iz}}=\frac{1}{\Gamma\left(1-\frac{iz}{2\pi}\right)}\,\exp\left[-\frac{iz}{2\pi}\log(-iz)\right],$ (211)
$G_{-}(z)$ $=\frac{K_{-}(z)}{\sqrt{iz}}=\frac{1}{\Gamma\left(1+\frac{iz}{2\pi}\right)}\,\exp\left[\frac{iz}{2\pi}\log(iz)\right]\,.$
Note that
@@ Appendix D Preliminary numerical extraction of perturbative coefficients
We solve the integral equation at $M=60$ values $Q_{j}\in[20,\,500]$ using $N(Q)=10Q+400$ quadrature points, subtract the known leading terms, and fit to the truncated ansatz
$R_{0}(Q)=\sum_{n=1}^{N_{\max}}\frac{a_{n0}+a_{n1}\log Q}{Q^{n}}$ (213)
by least squares, progressively increasing $N_{\max}$.
The design matrix has condition number growing with $N_{\max}$: only the first ten SVD modes carry signal, corresponding to five reliable coefficient pairs $(a_{n0},a_{n1})$. The stable coefficients ($n=1,\ldots,5$) exhibit an alternating sign pattern (+,,+,,+)(+,-,+,-,+), consistent with a Borel singularity on the positive real axis.
@@
The ratio test $r_{n}/n=a_{(n+1)0}/(n\,a_{n0})$ should converge to $1/(2\pi)\approx 0.159$. The computed ratios (using $n\leq 5$) give $r_{n}/n\approx 0.6$$1.0$, still far from the predicted value—this slow convergence is expected since the asymptotic regime requires $n\gg 2\pi$, and with only five reliable coefficients we are probing the pre-asymptotic regime.
A more direct test is to subtract the $N$-term perturbative approximation and look for the exponential tail. However, a quantitative analysis reveals that clean detection of the instanton is fundamentally impossible in double precision: at any $Q$, the perturbative remainder dominates $e^{-2\pi Q}$ by 9 or more orders of magnitude. The machine-epsilon threshold $e^{-2\pi Q}\approx 10^{-16}$ occurs at $Q\approx 5.9$, but at such a value the perturbative remainder after subtracting four terms is still $\sim\mathcal{O}(1)$. Extending to 15–20 reliable coefficients via extended-precision arithmetic would open a detection window.